OpenStudy (anonymous):

Evaluate the following expressions, when a=-2, b=3, c=-3: 2a(b

OpenStudy (anonymous):

\[2a(b ^{2} - a)\div c\]

OpenStudy (anonymous):

Please give working thx

OpenStudy (anonymous):

2a=2x-2=-4 b^2-a=3^2-(-2)=9+2=11 (-4)(11)/(-3)=44/3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

I have another one

OpenStudy (anonymous):

\[2a+ b \div a - c\]

OpenStudy (anonymous):

here u should follow priorities division before addition and subtraction -4+3/(-2)-(-3) = -4-1.5+3=-2.5

OpenStudy (anonymous):

not good

OpenStudy (anonymous):

-4+3 = -1

OpenStudy (anonymous):

and -2+3 = 1

OpenStudy (anonymous):

-1 div 1 = -1

OpenStudy (anonymous):

where im wrong people